It takes a person the same time to drive 150 miles as it takes a plane to fly 1350 miles. If the plane is flying 400 mph faster than the car, how fast is each traveling?
Answer by checkley77(12844)

D=RT OR T=D/R
150/X=1350(X+400)
1350X=150(X+400)
1350X=150X+60000
1350X-150X=60000
1200X=60000
X=60000/1200
X=50 MPH. FOR THE CAR.
50+400=450 MPH FOR THE PLANE.
PROOF:
150/50=1350/450
3=3